OpenStudy (anonymous):

Which of the following is equivalent to the statement below? If 3x + 1 = 10, then x ≠ 2.

OpenStudy (jiteshmeghwal9):

3x=10-1 3x=9 x=9/3=3

OpenStudy (anonymous):

If x ≠ 2, then 3x + 1 = 10. If 3x + 1 = 10, then x = 2. If x = 2, then 3x + 1 = 10. If x = 2, then 3x + 1 ≠ 10.

OpenStudy (jiteshmeghwal9):

If x = 2, then 3x + 1 ≠ 10